# Stochastic Thermodynamics of Non-reciprocally Interacting Particles and Fields

### v2

@@ A Energy Functional and Non-conservative Forces
The macroscopic Boltzmann weight $\mu_{i}$ of $\rho_{i}$ is identified using LDB [86]. It is composed of the ideal (Boltzmann entropic) and reciprocal interaction (energetic) parts denoted by $\mu_{i}^{id}=\ln{\rho_{i}}$ and $\mu_{i}^{int}=\sum_{j}{V}_{ij}\rho_{j}$ respectively, such that $\mu_{i}=\mu_{i}^{id}+\mu_{i}^{int}$. Where $V_{ij}=\Omega\left(e^{\beta v_{ij}^{r}+\beta v^{nr}_{ij}}-1\right)$ quantifies the contribution to $\mu_{i}$ due to $\rho_{j}$ attributed to the interaction between $\rho_{i}$ and $\rho_{j}$. $V_{ij}$ is the second Virial co-efficient for the interacting non-reciprocal fields. $\mu_{i}^{int}<0$ signifies attractive interactions experienced by $\rho_{i}$ due to other macrostates. Analogously, $\mu_{i}^{int}>0$ means a repulsive interactions experienced by $\rho_{i}$. $\mu_{i}$ is decomposed into its reciprocal and non-reciprocal Boltzmann weights:
$\begin{split}\mu_{i}^{r}=\ln{\rho_{i}}+\sum_{j}{V}_{ij}^{r}\rho_{j},\hskip 14.22636pt{F}^{nr}_{i}=\sum_{j}{V}_{ij}^{nr}\rho_{j}.\end{split}$ (24)
The reciprocal and non-reciprocal interaction coefficients for the $\rho_{i}$ due to $\rho_{j}$ are denoted by $V_{ij}^{r}$ and $V_{ij}^{nr}$ respectively.
@@
Importantly, ${V}_{ij}^{r}={V}_{ji}^{r}$ and ${V}_{ij}^{nr}=-{V}_{ji}^{nr}$ is satisfied due to ${v}_{ij}^{r}={v}_{ji}^{r}$ and ${v}_{ij}^{nr}=-{v}_{ji}^{nr}$. Hence, the microscopic ‘orthogonal gauge fixing’ introduced in section 2.1 is scale-invariant. This ensures the uniqueness of reciprocal non-reciprocal decomposition on the coarse-grained macroscale and its exact mapping to the microscopic counterpart. Importantly, the thermodynamic quantities are also scale-invariant, ensuring the consistency between the microscopic and macroscopic descriptions. Importantly, other gauges do not respect the scale invariance and lead to physically incorrect decomposition of EPR.
The non-linear dependence of the interaction coefficients ${V}_{ij}^{r}$ and ${V}_{ji}^{nr}$ are attributed to incorporating the Posissonian statistics of the microscopic occupancy variable [86]. $\mu_{i}^{r}$ is derived from a unique macroscopic global energy functional $E$ defined for the system such that $\mu_{i}^{r}=\delta E/\delta\rho_{i}$.
@@
$E$ quantifies the total reciprocal interaction energy $E^{int}$ between macrostates and the macroscopic Boltzmann entropic term $S^{b}$, such that $E=E^{int}-S^{b}$. The macroscopic non-conservative driving forces along the reactive and diffusive transitions are given by it’s microscopic counterparts, $F^{ch}_{\gamma\gamma^{\prime}}=f_{\gamma\gamma^{\prime}}^{ch}$ and $\vec{F}^{sp}_{i}=l\vec{f}_{i}^{sp}$
@@ B Equilibrium energy functional
$E$ is not the equilibrium energy functional $E^{eq}$. It is trivially verified by using $v_{ij}^{nr}=0,\vec{f}_{i}^{sp}=0,f_{\gamma\gamma^{\prime}}^{ch}=0$ leads to $E^{eq}=E(V_{ij}^{eq})$, where, ${V}_{ij}^{eq}=\Omega\left[e^{\beta v_{ij}^{r}}-1\right]$ is the equilibrium second Virial coefficient. Nonetheless, $E^{eq}$ gives a lower bound on $E$, $|{E}|\geq|{E}^{eq}|$ due to $|{V}_{ij}^{r}|\geq|{V}_{ij}^{eq}|$. The equality hold for $v_{ij}^{nr}=0,\forall\>i,j$. In contrast to $E^{eq}$, $E$ is defined for non-reciprocal systems. In addition, it incorporates the effect of non-reciprocal interactions on the symmetric macrostate correlations through $\cosh{\left(\beta v^{nr}_{ij}\right)}$ term in eq. 25. Thus, $E$ is a Lyapunov functional for non-reciprocal systems in the absence of external driving forces and vanishing non-reciprocal vorticity currents.
@@ A Local Detailed Balance
The macroscopic reactive and diffusive transitions are defined as $\Delta_{\gamma\gamma^{\prime}}:\rho_{\gamma^{\prime}}\to\rho_{\gamma}$ and $\Delta_{i}^{\vec{\mathcal{D}}}:\rho_{i}\to\rho_{i}^{\vec{\mathcal{D}}}$ respectively. ${\vec{{\mathcal{D}}}}$ quantifies the direction vector for the diffusive transition. We identify the LDB using the macroscopic transition probability measure using the DPFT [86]. The LDB for the reactive ($\Delta_{\gamma\gamma^{\prime}}$) and diffusive (${\Delta}_{i}^{\vec{\mathcal{D}}}$) macroscopic transitions read:
[…]
The LDB constrains the ratio of the forward and backward transitions (a dynamic quantity) using $\mu_{i}$ (a thermodynamic quantity). For $\Delta_{\gamma\gamma^{\prime}}$, we define the macroscopic affinity $A_{\gamma\gamma^{\prime}}=\mu_{\gamma^{\prime}}-\mu_{\gamma}+F_{\gamma\gamma^{\prime}}$ and it’s symmetric counterpart $U_{\gamma\gamma^{\prime}}=\mu_{\gamma^{\prime}}+\mu_{\gamma}$. Similarly, for ${\Delta}_{i}^{\vec{\mathcal{D}}}$, we define ${A}_{i}^{\vec{\mathcal{D}}}=\mu_{i}-\mu_{i}^{\vec{\mathcal{D}}}+\vec{\mathcal{D}}\cdot\vec{F}_{i}^{sp}$ and ${U}_{i}^{\vec{\mathcal{D}}}=2\mu_{i}$. Macroscopic and microscopic self-propulsion are related by, $\vec{F}_{i}^{sp}=l\vec{f}_{i}^{sp}$.
@@ B Generalized Macroscopic Fluctuating Dynamics
The deterministic evolution of $\rho_{i}$ is given by the ‘most likelihood path’ of the Doi-Peliti action, namely ‘Instanton’. The ‘dominant macrostate fluctuations’ are characterized by the local curvature of the ‘Instanton’ [86]. The stochastic equation of motion for $\rho_{i}$ is [86]:
@@
We choose the convention as the outward transition currents from $\rho_{i}$ ($-J_{\gamma\gamma^{\prime}}$ and $-\nabla\cdot\vec{J}_{i}^{\vec{\mathcal{D}}}$), represented by a negative sign. $\hat{\xi}_{\gamma\gamma^{\prime}}^{\mathcal{R}}$ and $\hat{\xi}_{i}^{\mathcal{D}}$ are standard Gaussian white noise with unit variance and vanishing mean. $\{\Delta_{\gamma\gamma^{\prime}}\}$ denotes the set of all macroscopic reactive transitions. The reactive and diffusive transition currents are:
@@
Where $\nabla^{\vec{\mathcal{D}}}$ and $\tilde{d}_{i}^{\vec{{\mathcal{D}}}}=d_{i}l^{2}\cosh{(\vec{\mathcal{D}}\cdot\vec{F}_{i}^{sp}/2)}$ are the gradient and the macroscopic diffusion constant along $\vec{\mathcal{D}}$ respectively. Similarly, the Laplacian along $\vec{\mathcal{D}}$ is denoted by $\Delta^{\vec{\mathcal{D}}}$. Here, $l$ is the microscopic diffusive length scale. $\vec{J}_{i}^{sp}=2d_{i}le^{\mu_{i}}\sinh{(\vec{\mathcal{D}}\cdot\vec{F}_{i}^{sp}/2)}$ gives the macroscopic self-propulsion current. Choosing the basis vector $\vec{\mathcal{D}}=\{\parallel,\perp\}$ as parallel and perpendicular to $\vec{F}_{i}^{sp}$, the simplified form reads $-\nabla\cdot\vec{J}_{i}^{\vec{\mathcal{D}}}=\tilde{d}_{i}^{\parallel}\Delta^{\parallel}e^{\mu_{i}}+d_{i}^{\perp}\Delta^{\perp}e^{\mu_{i}}+\nabla^{\parallel}\cdot\vec{J}_{i}^{sp}$. Here, $d_{i}^{\parallel}=d_{i}l^{2}\cosh{(F_{i}^{sp}/2)}$ and $d_{i}^{\perp}=d_{i}l^{2}$ reveal that self-propulsion renormalizes the diffusion coefficients differently in the parallel and orthogonal directions to the self-propulsion driving force. Without loss of generality, we consider the scaled macroscopic diffusion constants $\tilde{d}_{i}=d_{i}l^{2}$ or simply fix $l=1$ equivalent to using the diffusive length-scale a unit of spacial distance. The continuum limit $l\to 0$ leads to $\tilde{d}_{i}^{\parallel}=\tilde{d}_{i}^{\perp}=\lim_{l\to 0}d_{i}l^{2}$ and $J_{i}^{sp}=\tilde{d}_{i}^{\parallel}e^{\mu_{i}}f_{i}^{sp}$ [86]. Hence, the macrostates lose the directional dependence (anisotropy) of diffusion coefficients due to oversimplification (coarse-graining) of the diffusive length scale. However, the anisotropic diffusion coefficients have led to the formation of novel phases [125, 126, 127, 128, 129], highlighting its importance.
We define the mobility for the diffusive and reactive transitions as follows.
@@
It characterizes the strength of the macroscopic transition currents generated by the macroscopic transition affinities. It plays the same role as diffusive mobility and transport coefficients for $\vec{J}_{i}^{\vec{\mathcal{D}}}$ and $\vec{J}_{i}^{sp}$ respectively. For repulsive interactions experienced by $\rho_{i}$ (that is, $\mu_{i}>0$), the amplitude of the transition mobilities increases exponentially, signifying escape from a thermodynamically unfavorable state. Similarly, the transition currents are exponentially suppressed for $\rho_{i}$ experiencing the attractive interaction (that is, $\mu_{i}<0$). The variances of the reactive and diffusive currents read [86]:
@@
The eq. 31 is the Einstein-Smoluchowski relation that connects current fluctuations to transition mobility, where $\Omega$ plays the role of inverse temperature $\beta$ [124, 86]. Importantly, the same mobility controls the mean current in eq. 29 through a hyperbolic relation. Consequently, this hyperbolic relationship leads to the fluctuation response relation. We define traffic as the sum of the modulus of forward and backward currents [130, 131, 132, 133, 134]. The traffic for the reactive and diffusive transitions is:
[…]
Traffic in eq. 32 is related to the scaled variance of the transition currents eq. 31, $T_{\gamma\gamma^{\prime}}=2\Omega B_{\gamma\gamma^{\prime}}\left(\{\rho\}\right)$ and $T_{i}^{\vec{\mathcal{D}}}=2\Omega B_{i}^{\vec{\mathcal{D}}}\left(\{\rho\}\right)$.
Although eq. 28 is represented in the continuum macroscopic limit $l\to 0$, equivalent analogues exist for systems with finite $l$, for example, lattice gas models $l=1$. The discrete analogue of the Laplacian and gradient operators should replace the continuum space counterpart, similar to the reactive transition dynamics [86].

### v3

@@ A Energy Functional and Non-conservative Forces
The macroscopic Boltzmann weight $\mu_{i}$ of the macrostate $\rho_{i}$ is identified using the LDB [87]. It consists of an ideal (entropic) part $\mu_{i}^{id}=\ln{\rho_{i}}$ and a reciprocal interaction (energetic) part $\mu_{i}^{int}=\sum_{j}{V}_{ij}\rho_{j}$, such that $\mu_{i}=\mu_{i}^{id}+\mu_{i}^{int}$. Here, $V_{ij}=\Omega\left(e^{\beta v_{ij}^{r}+\beta v^{nr}_{ij}}-1\right)$ quantifies the contribution to $\mu_{i}$ due to the presence of $\rho_{j}$ arising from the interaction between $\rho_{i}$ and $\rho_{j}$. $V_{ij}$ serves as the second Virial coefficient for the interacting non-reciprocal fields. $\mu_{i}^{int}<0$ indicates attractive interactions experienced by $\rho_{i}$, whereas $\mu_{i}^{int}>0$ corresponds to repulsive interactions. Analogously, $\mu_{i}$ can be decomposed into reciprocal and non-reciprocal macroscopic Boltzmann weights:
$\begin{split}\mu_{i}^{r}&=\ln{\rho_{i}}+\sum_{j}{V}_{ij}^{r}\rho_{j},\\ {F}^{nr}_{i}&=\sum_{j}{V}_{ij}^{nr}\rho_{j},\end{split}$ (24)
where $V_{ij}^{r}$ and $V_{ij}^{nr}$ denote the reciprocal and non-reciprocal interaction coefficients for $\rho_{i}$ due to $\rho_{j}$, respectively:
@@
Table 2 summarizes the notation used for macroscopic interacting fields. Importantly, $V_{ij}^{r}=V_{ji}^{r}$ and $V_{ij}^{nr}=-V_{ji}^{nr}$, as required by $v_{ij}^{r}=v_{ji}^{r}$ and $v_{ij}^{nr}=-v_{ji}^{nr}$. This ensures that the microscopic ‘orthogonal gauge fixing’ is scale-invariant, guaranteeing a unique and exact decomposition of reciprocal and non-reciprocal contributions at the coarse-grained macroscopic level. Alternative gauge choices would violate scale invariance and lead to physically incorrect decompositions of the EPR.
Table 2: Notations for macroscopic interacting fields. The subscript $i$ denotes the type of the macrostate $\rho_{i}$. For interaction coefficients between two fields, two indices are used, following the ‘left-right’ convention (‘influenced-influencer’). For macroscopic transitions, the initial-to-final (‘right-to-left’) convention is used.
Notation Significance
$V_{ij}$ The macroscopic interaction potential energy experienced by the macrostate $\rho_{i}$ due to the presence of $\rho_{j}$.
$V_{ij}^{r}$ The reciprocal component of $V_{ij}$ that satisfies the macroscopic ‘actio=reactio’ symmetry $V_{ij}^{r}=V_{ji}^{r}$.
$V_{ij}^{nr}$ The non-reciprocal component of $V_{ij}$ that satisfies the macroscopic ‘actio=-reactio’ anti-symmetry $V_{ij}^{nr}=-V_{ji}^{nr}$.
$\mu_{i}$ The macroscopic Boltzmann weight of the macrostate $\rho_{i}$, quantifying the total thermodynamic cost of placing a macrostate $\rho_{i}$.
$\mu_{i}^{r}$ The reciprocal part of $\mu_{i}$ that respects the macroscopic ‘actio=reactio’ symmetry.
$F^{nr}_{i}$ The non-reciprocal part of $\mu_{i}$ that violates the macroscopic ‘actio=reactio’ symmetry.
$E^{int}$ The total macroscopic reciprocal interaction energy for the system.
$\Delta_{\gamma\gamma^{\prime}}$ A reactive transition for macrostate $\rho_{\gamma^{\prime}}$ to change its chemical state to macrostate $\rho_{\gamma}$: $\rho_{\gamma^{\prime}}\to\rho_{\gamma}$.
$F_{\gamma\gamma^{\prime}}^{ch}$ An external ‘chemical’ driving force along the transition $\Delta_{\gamma\gamma^{\prime}}$, supported by an enzymatic thermodynamic reservoir.
$K_{\gamma\gamma^{\prime}}$ The reactive transition rate corresponding to $\Delta_{\gamma\gamma^{\prime}}$.
$\Delta_{i}^{\vec{\mathcal{D}}}$ A diffusive transition for macrostate $\rho_{i}$ in the direction $\vec{\mathcal{D}}$: $\rho_{i}\to\rho_{i}^{\vec{\mathcal{D}}}$.
$\vec{F}_{i}^{sp}$ A ‘self-propulsion’ driving force along the transition $\Delta_{i}^{\vec{\mathcal{D}}}$, supported by a colloidal reservoir around the field.
$K_{i}^{\vec{\mathcal{D}}}$ The diffusive transition rate for $\Delta_{i}^{\vec{\mathcal{D}}}$.
The non-linear dependence of $V_{ij}^{r}$ and $V_{ij}^{nr}$ arises from incorporating the Poissonian statistics of the microscopic occupancy variables [87]. The reciprocal macroscopic weight $\mu_{i}^{r}$ is derived from the global macroscopic energy functional $E$, defined such that $\mu_{i}^{r}=\delta E/\delta\rho_{i}$:
@@
$E$ quantifies the total reciprocal interaction energy $E^{int}$ and the macroscopic Boltzmann entropic term $S^{b}$, such that $E=E^{int}-S^{b}$. Consequently, $\mu_{i}=\mu_{i}^{r}+F_{i}^{nr}$ represents a macroscopic Helmholtz-Hodge decomposition of $\mu_{i}$ into symmetry-preserving ($\mu_{i}^{r}$) and symmetry-violating ($F_{i}^{nr}$) components. The macroscopic non-conservative driving forces along reactive and diffusive transitions are related to their microscopic counterparts as $F^{ch}_{\gamma\gamma^{\prime}}=f_{\gamma\gamma^{\prime}}^{ch}$ and $\vec{F}^{sp}_{i}=l\vec{f}_{i}^{sp}$.
@@ B Equilibrium energy functional
The functional $E$ is not the equilibrium energy functional $E^{eq}$. Setting $v_{ij}^{nr}=0$, $\vec{f}_{i}^{sp}=0$, and $f_{\gamma\gamma^{\prime}}^{ch}=0$ yields $E^{eq}=E(V_{ij}^{eq})$, where $V_{ij}^{eq}=\Omega\left[e^{\beta v_{ij}^{r}}-1\right]$ is the second Virial coefficient for equilibrium reciprocal interactions. Notably, $E$ satisfies $|E|\geq|E^{eq}|$, with equality only if $v_{ij}^{nr}=0$ for all $i,j$. Unlike $E^{eq}$, the functional $E$ incorporates non-reciprocal interactions and their effect on symmetric macrostate correlations through the $\cosh(\beta v_{ij}^{nr})$ term in eq. 25. Therefore, $E$ serves as a Lyapunov functional for non-reciprocal systems in the absence of external driving forces and vanishing non-reciprocal vorticity currents.
@@ A Local Detailed Balance
The macroscopic reactive and diffusive transitions are defined as $\Delta_{\gamma\gamma^{\prime}}:\rho_{\gamma^{\prime}}\to\rho_{\gamma}$ and $\Delta_{i}^{\vec{\mathcal{D}}}:\rho_{i}\to\rho_{i}^{\vec{\mathcal{D}}}$, respectively, where $\vec{{\mathcal{D}}}$} quantifies the direction vector for the diffusive transition. Using the macroscopic transition probability measure from the DPFT [87], the macroscopic transition rates for the reactive ($\Delta_{\gamma\gamma^{\prime}}$) and diffusive (${\Delta}_{i}^{\vec{\mathcal{D}}}$) transitions are:
$\begin{split}K_{\gamma\gamma^{\prime}}&=d_{\gamma\gamma^{\prime}}e^{\mu_{\gamma^{\prime}}+\frac{1}{2}F_{\gamma\gamma^{\prime}}^{ch}},\hskip 14.22636ptK_{i}^{\vec{\mathcal{D}}}=\tilde{d}_{i}^{\vec{\mathcal{D}}}e^{\mu_{i}+\frac{1}{2}\vec{\mathcal{D}}\cdot\vec{F}_{i}^{sp}}.\end{split}$ (27)
Here, $\tilde{d}_{i}^{\vec{\mathcal{D}}}=d_{i}l^{2}\cosh{(\vec{\mathcal{D}}\cdot\vec{F}_{i}^{sp}/2)}$ is the macroscopic diffusion constant along $\vec{\mathcal{D}}$, measured using the microscopic diffusive length scale $l$. The local detailed balance (LDB) constructed from eq. 27 for the reactive ($\Delta_{\gamma\gamma^{\prime}}$) and diffusive (${\Delta}_{i}^{\vec{\mathcal{D}}}$) transitions reads [87]:
$\begin{split}\frac{K_{\gamma\gamma^{\prime}}}{K_{\gamma^{\prime}\gamma}}&=e^{\mu_{\gamma^{\prime}}-\mu_{\gamma}+F_{\gamma\gamma^{\prime}}^{ch}},\hskip 14.22636pt\frac{K_{i}^{\vec{\mathcal{D}}}}{K_{i}^{(\vec{\mathcal{D}})^{-1}}}=e^{\mu_{i}-\mu_{i}^{\vec{\mathcal{D}}}+\vec{\mathcal{D}}\cdot\vec{F}_{i}^{sp}}.\end{split}$ (28)
The LDB constrains the ratio of forward and backward transitions (a dynamic quantity) using $\mu_{i}$ (a thermodynamic quantity). For $\Delta_{\gamma\gamma^{\prime}}$, we define the macroscopic affinity $A_{\gamma\gamma^{\prime}}=\mu_{\gamma^{\prime}}-\mu_{\gamma}+F_{\gamma\gamma^{\prime}}$ and its symmetric counterpart $U_{\gamma\gamma^{\prime}}=\mu_{\gamma^{\prime}}+\mu_{\gamma}$. Similarly, for ${\Delta}_{i}^{\vec{\mathcal{D}}}$, we define ${A}_{i}^{\vec{\mathcal{D}}}=\mu_{i}-\mu_{i}^{\vec{\mathcal{D}}}+\vec{\mathcal{D}}\cdot\vec{F}_{i}^{sp}$ and ${U}_{i}^{\vec{\mathcal{D}}}=2\mu_{i}$. Macroscopic and microscopic self-propulsion are related by $\vec{F}_{i}^{sp}=l\vec{f}_{i}^{sp}$. Equations 27 and 28 are the macroscopic analogs of the microscopic eqs. 3 and 4, respectively. Importantly, eq. 27 provides a more fundamental formulation than eq. 28, as it does not assume a timescale separation between different transition rates. Since eq. 28 omits the time-symmetric part of the transition rates by taking the ratio of forward and backward transitions, it is analogous to the microscopic counterpart discussed previously.
@@ B Generalized Macroscopic Fluctuating Dynamics
The deterministic evolution of $\rho_{i}$ is given by the ‘most likelyhood path’ of the Doi-Peliti action, namely the ‘Instanton’. The ‘dominant macrostate fluctuations’ are characterized by the local curvature of the ‘Instanton’ [87]. The stochastic equation of motion for $\rho_{i}$ is [87]:
@@
We adopt the convention that outward transition currents from $\rho_{i}$ ($J_{\gamma\gamma^{\prime}}$ and $\nabla\cdot\vec{J}_{i}^{\vec{\mathcal{D}}}$) are represented with a negative sign. $\hat{\xi}_{\gamma\gamma^{\prime}}^{\mathcal{R}}$ and $\hat{\xi}_{i}^{\mathcal{D}}$ are standard Gaussian white noises with unit variance and zero mean. The set $\{\Delta_{\gamma\gamma^{\prime}}\}$ denotes all macroscopic reactive transitions. The reactive and diffusive transition currents are:
@@
Here, $\nabla^{\vec{\mathcal{D}}}$ is the gradient along $\vec{\mathcal{D}}$, and similarly, the Laplacian along $\vec{\mathcal{D}}$ is denoted by $\Delta^{\vec{\mathcal{D}}}$. The macroscopic self-propulsion current is $\vec{J}_{i}^{sp}=2d_{i}le^{\mu_{i}}\sinh{(\vec{\mathcal{D}}\cdot\vec{F}_{i}^{sp}/2)}$. Choosing the basis vector $\vec{\mathcal{D}}=\{\parallel,\perp\}$ as parallel and perpendicular to $\vec{F}_{i}^{sp}$, the simplified form reads $\nabla\cdot\vec{J}_{i}^{\vec{\mathcal{D}}}=\tilde{d}_{i}^{\parallel}\Delta^{\parallel}e^{\mu_{i}}+d_{i}^{\perp}\Delta^{\perp}e^{\mu_{i}}+\nabla^{\parallel}\cdot\vec{J}_{i}^{sp}$. Here, $d_{i}^{\parallel}=d_{i}l^{2}\cosh{(F_{i}^{sp}/2)}$ and $d_{i}^{\perp}=d_{i}l^{2}$, indicating that self-propulsion renormalizes the diffusion coefficients differently in directions parallel and perpendicular to the self-propulsion force. Without loss of generality, we consider the scaled macroscopic diffusion constants $\tilde{d}_{i}=d_{i}l^{2}$ or equivalently fix $l=1$, using the diffusive length scale as the unit of spatial distance. In the continuum limit $l\to 0$, $\tilde{d}_{i}^{\parallel}=\tilde{d}_{i}^{\perp}=\lim_{l\to 0}d_{i}l^{2}$ and $J_{i}^{sp}=\tilde{d}_{i}^{\parallel}e^{\mu_{i}}f_{i}^{sp}$ [87], so the macrostates lose the directional dependence of the diffusion coefficients. The anisotropic diffusion coefficients, however, have been shown to generate novel phases [131, 132, 133, 134, 135], highlighting their importance.
We define the mobility for diffusive and reactive transitions as:
@@
Mobilities characterize the strength of macroscopic transition currents generated by the macroscopic transition affinities. They play the same role as diffusive mobility and transport coefficients for $\vec{J}_{i}^{\vec{\mathcal{D}}}$ and $\vec{J}_{i}^{sp}$, respectively. For repulsive interactions experienced by $\rho_{i}$ (i.e., $\mu_{i}>0$), the amplitude of the transition mobilities increases exponentially, signifying escape from a thermodynamically unfavorable state. Similarly, the transition currents are exponentially suppressed for $\rho_{i}$ experiencing attractive interactions ($\mu_{i}<0$). The variances of the reactive and diffusive currents are [87]:
@@
Equation 32 is the Einstein-Smoluchowski relation connecting current fluctuations to transition mobility, where $\Omega$ plays a role analogous to the inverse temperature $\beta$ [130, 87]. Importantly, the same mobility controls the mean current in eq. 30 through a hyperbolic relation, which gives rise to the fluctuation-response relation.
We define the traffic as the sum of the magnitudes of the forward and backward currents [136, 137, 138, 139, 140]. For reactive and diffusive transitions, the traffic reads:
$\begin{split}T_{\gamma\gamma^{\prime}}=2D_{\gamma\gamma^{\prime}}\left(\{\rho\}\right)\cosh{\left(\frac{A_{\gamma\gamma^{\prime}}}{2}\right)},\hskip 19.91684ptT_{i}^{\vec{\mathcal{D}}}=2D_{i}\left(\{\rho\}\right).\end{split}$ (33)
Traffic in eq. 33 is related to the scaled variance of the transition currents in eq. 32: $T_{\gamma\gamma^{\prime}}=2\Omega B_{\gamma\gamma^{\prime}}\left(\{\rho\}\right)$ and $T_{i}^{\vec{\mathcal{D}}}=2\Omega B_{i}^{\vec{\mathcal{D}}}\left(\{\rho\}\right)$.
The mean current and traffic (eqs. 30 and 33) exhibit $\sinh()$ and $\cosh()$ forms due to their thermodynamically consistent exponential parametrization in the affinity $A$. However, they have a fundamental thermodynamic meaning independent of exponential parametrization: namely, the mean current and traffic correspond to the time-antisymmetric and time-symmetric components of the non-equilibrium currents [89, 90]. Consequently, they characterize the thermodynamic length and the dynamical activity for the transition, which physically correspond to the total directional transition current and the inverse timescale for the transition [89, 90]. The shortcomings of the Glauber-dynamics and Metropolis-dynamics discussed in section 2.2 are better understood with eq. 29, where different transitions contribute with different ‘diffusivity’ or ‘variance’, quantified by the traffic. This exactly preserves the different timescales associated with different transitions that contribute to the macrostate $\rho_{i}$. Therefore, due to the thermodynamically consistent exact coarse-graining procedure [87], eq. 29 serves as a foundational cornerstone for future studies to investigate macroscopic stochastic thermodynamics without strong assumptions on timescale separation.
Although eq. 29 is written in the continuum macroscopic limit $l\to 0$, equivalent analogs exist for systems with finite $l$, e.g., lattice gas models with $l=1$. In such cases, the discrete analogs of the Laplacian and gradient operators replace the continuum-space counterparts, similar to reactive transition dynamics [87].